### v2

# Control of active field theories at minimal dissipation

Abstract
Advances in experimental techniques enable the precise manipulation of a large variety of active systems, which constantly dissipate energy to sustain nonequilibrium phenomena without any equilibrium equivalent. To design novel materials out of active systems, an outstanding challenge is to rationalize how material properties can be optimally controlled by applying external perturbations. However, equilibrium thermodynamics is inadequate to guide the control of such nonequilibrium systems. Therefore, there is a dire need for a novel framework to provide a systematic toolbox for the thermodynamic control of active matter. Here, we build an optimization procedure for generic active field theories within a thermodynamically consistent formulation. Central to our approach is the distinction between the protocol heat, which is dissipated only during manipulation, and the total heat, which also accounts for the post-manipulation dissipation. We demonstrate that the latter generically features a global minimum with respect to the protocol duration. We deploy our versatile approach to an active theory of phase separation, and examine the scalings of the optimal protocol duration with respect to activity and system size. Remarkably, we reveal that the landscape of steady-state dissipation regulates the crossover between optimal control strategies for a finite duration.
Introduction
Equilibrium thermodynamics states that manipulating passive systems at minimal dissipation can only be achieved in the quasistatic regime of infinitely slow perturbation [1]. Active systems evade such a constraint due to the constant consumption of energy at the basis of their internal sustained motion and/or mechanical stress [2, 3]. Then, the dissipation resulting from manipulating active systems can exhibit a minimum at a finite protocol duration [4]. This behaviour defies conventional intuition, and requires a novel optimization approach beyond that of passive systems.
Experimental techniques have demonstrated the ability to reliably manipulate active systems; for instance, using magnetic fields [5, 6], light sources [7, 8, 9, 10, 11], or nematic pumps [12]. These developments call for a theoretical framework to guide experiments from empirical to optimal control. The methods of stochastic thermodynamics set the basis for building such a framework [13]. These methods have mostly been deployed in systems with a few degrees of freedom [14, 15, 16], although some studies have considered controlling equilibrium spin models [17, 18, 19]. To control active systems [20], previous theoretical studies have offered some strategies for the navigation [21, 22, 23, 24] and confinement [25, 26, 27, 28] of particles, some of which using machine learning [29, 30]. Other studies have addressed the control of hydrodynamic theories describing active matter as continuum materials [31, 32, 33, 34, 35]. Remarkably, most control strategies put forward so far at the hydrodynamic level neglect the role of noise, which hinders their use in active systems where fluctuations are predominant.
The collective behaviors in passive systems have been successfully studied using hydrodynamic field theories [36]. Similarly, nonequilibrium field theories provide a broad framework to describe the large-scale features of active systems [2, 37, 38], such as motility-induced phase separation [39], collective directed motion [40], and odd viscosity [41, 42, 43]. In particular, field theories delineate distinct classes of systems based on the symmetries of their microscopic interaction. Most of these theories are built to capture some nonequilibrium collective behaviours without any energetic consideration. Instead, recent studies have proposed thermodynamically consistent formulations which entail unambiguous quantification of dissipation [44, 45, 38, 46, 47, 48, 49, 50] by describing how active systems couple with equilibrium reservoirs [13, 51, 52, 53].
Figure 1: (a) During manipulation ($0<t<\tau$), the system changes its state under the protocol $a(t)$: the system accumulates external work ($W_{\rm ext}$), and dissipates protocol heat ($Q_{\rm p}$) into the thermostat. During relaxation ($\tau<t<\tau+\tau_{\rm r}$), the system changes state at constant $a(t)=a_{\tau}$: there is no external work, and the system dissipates relaxation heat ($Q_{\rm r}$). The active work $W_{\rm act}>0$ is produced during both manipulation and relaxation. (b) The protocol heat $Q_{\rm p}$ vanishes for instantaneous manipulation ($\tau=0$), and can feature a local maximum at a finite duration ($\tau=\tau_{\rm max}$). For passive systems, $Q_{\rm p}$ saturates at large $\tau$; for active systems, $Q_{\rm p}$ diverges at large $\tau$ due to internal activity, and can feature a local minimum ($\tau=\tau_{\rm p}$). (c) The total heat $Q_{\rm t}=Q_{\rm p}+Q_{\rm r}$ has a non-zero value at $\tau=0$. For passive systems, $Q_{\rm t}$ decreases towards a finite value; for active systems, $Q_{\rm t}$ diverges at large $\tau$, and always features a local minimum ($\tau=\tau_{\rm t}$).
In this paper, we search for generic control strategies in field theories of active matter. We propose a versatile framework to predict the optimal protocols which change material properties by tuning some experimentally accessible parameters. Inspired by a previous optimization procedure, originally proposed for passive systems [54] and recently adapted to active matter [4], we focus on minimizing the heat generated by smooth and slow protocols. In contrast with [54, 4], we argue that the protocol heat $Q_{\rm p}$, which is dissipated only during the protocol, is less adequate for optimization than the total heat $Q_{\rm t}$, which also accounts for the post-protocol dissipation [Fig. 1(a)]. For passive systems, $Q_{\rm t}$ is optimal at long protocol duration, as expected from standard thermodynamics [1], while it reaches a minimum at a finite duration for active systems [Fig. 1(b)]. Instead, $Q_{\rm p}$ is always zero for a vanishing protocol duration, and need not feature any other local minimum even for active systems [Fig. 1(c)].
We apply our framework to a thermodynamically consistent field theory, referred to as Chemical Model B (CMB), which describes nonequilibrium phase separation. Using dynamical response theory, we provide analytical predictions for protocols driving some homogeneous configurations of the system, and examine the scalings of the optimal duration with respect to system size and activity. Remarkably, we reveal a crossover between control strategies, and rationalize it based on the landscape of steady-state dissipation. Overall, our results demonstrate the potential of our novel approach to guiding control in a large class of active systems.
Results
The role of relaxation: Protocol vs total heat
We consider the external manipulation of a system at constant temperature $T$ through a finite-time protocol varying the parameter $a(t)$ from $a_{0}=a(0)$ to $a_{\tau}=a(\tau)$. The system is assumed to be initially ($t=0$) at steady state [Fig. 1(a)]. For any finite protocol duration $\tau$, the protocol is not quasi-static, so that the system is not in steady state at any time $t\in(0,\tau)$. After the protocol ($t>\tau$), the parameter $a(t>\tau)=a_{\tau}$ is constant, and the system relaxes towards the stationary configuration given by $a_{\tau}$. We define the post-protocol time $\tau_{\rm r}$, which is far larger than the relaxation of the slowest mode, such that the system reaches a steady state at time $\tau+\tau_{\rm r}$.
We denote by $Q_{\rm p}$ the protocol heat dissipated by the system during the protocol. During the subsequent relaxation, an additional contribution, referred to as the relaxation heat $Q_{\rm r}$ is dissipated, although the parameter $a(t>\tau)=a_{\tau}$ is constant. The total heat $Q_{\rm t}$, which is dissipated during the protocol and the post-protocol relaxation, follows as $Q_{\rm t}=Q_{\rm p}+Q_{\rm r}$. Clearly, $(Q_{\rm p},Q_{\rm r},Q_{\rm t})$ all strongly depend on the shape of the protocol $a(t)$ and its duration $\tau$. Conservation of energy states that $Q_{\rm t}$ is given by the sum of the total work $W_{\rm t}$ and the change in the system energy $E$:
$Q_{\rm t}=W_{\rm t}+E(t=0)-E(t={\tau+\tau_{\rm r}})\,,$ (1)
where $W_{\rm t}$ can be decomposed in terms of the contributions produced by the external manipulation ($W_{\rm ext}$, accumulated only during the protocol) and by any non-conservative active forces ($W_{\rm act}$, accumulated during both the protocol and the subsequent relaxation):
$W_{\rm t}=W_{\text{ext}}+W_{\text{act}}\,.$ (2)
In the absence of external manipulation, the quantity $W_{\text{act}}$ is the energetic cost needed to sustain non-equilibrium steady states [55]. In the presence of fluctuations, Eqs. (1) and (2) take the same form for the stochastic definitions of heat, work, and energy [13, 51]. In what follows, we focus on average values; in particular, the averages of $E$ at times $t=0$ and $t=\tau+\tau_{\rm r}$ are evaluated in steady state, and thus are independent of the protocol.
For passive systems ($W_{\text{act}}=0$), Eq. (1) reduces to the first law of thermodynamics in its standard form, so minimizing either $Q_{\rm t}$ or $W_{\rm ext}$ leads to the same optimal protocol. The second law of thermodynamics states that $W_{\rm ext}$ can only be larger than the difference between the Helmholtz free energy evaluated at $a_{0}$ and $a_{\tau}$ [1]. Likewise, $Q_{\rm t}$ is bounded from below by the difference of the system entropy at $a_{0}$ and $a_{\tau}$, and the bound is only saturated for quasi-static protocols. For instantaneous protocols ($\tau=0$), $Q_{\rm p}$ vanishes since there is no sufficient time to dissipate heat during the protocol [Fig. 1(b)], whereas $Q_{\rm t}$ has a non-zero value as a signature of the dissipation ensued during the post-protocol relaxation [Fig. 1(c)].
The situation is markedly different for active systems. In steady state, $W_{\text{act}}$ increases extensively with time to sustain the internal processes at the basis of activity. Then, at large duration $\tau$, both $Q_{\rm t}$ and $Q_{\rm p}$ linearly increase with $\tau$, in contrast with passive systems, whereas $W_{\rm ext}$ saturates as in the passive case. It follows that $Q_{\rm t}$ always features a global minimum at finite $\tau$ [Fig. 1(c)], which achieves the best trade-off between the dissipation coming from internal processes and that stemming from external manipulation. Remarkably, $Q_{\rm p}$ vanishes at $\tau=0$ as in passive systems, and it may feature a local minimum $\tau_{\rm p}$ under certain conditions [Fig. 1(b)], in contrast with passive systems. For instance, for a scalar field theory, these conditions require a small activity (and also small system sizes, if the field is conserved), as we discuss in detail in the next sections.
In this paper, minimizing $Q_{\rm t}$ (rather than $W_{\text{ext}}$ [56] or $Q_{\rm p}$ [4]) is the objective of our optimization. In particular, our aim is to quantify the optimal duration $\tau_{\rm t}$ at minimal $Q_{\rm t}$, which separates the two regimes of dissipation inherent to active systems [Fig. 1(b)]. Specifically, we deploy a systematic optimization framework [54, 4] in a thermodynamically consistent formulation of active field theories [38], examine the corresponding scalings of $\tau_{\rm t}$, and unveil a crossover between optimal protocols.
Energetics of active field theories
We describe the energy exchange between a given active system and its surrounding reservoirs within linear irreversible thermodynamics [57, 38]; the energy reservoir absorbs the heat dissipated by the system, and the internal active driving is sustained by some active driving reservoir (typically chemical fuel). In practice, we consider here the coupling of two scalar fields: the density of active particles $\phi({\bf r},t)$, and the concentration of fuel molecules $n({\bf r},t)$. The specific nature of active driving is unimportant  [^1] What is important is the time-reversal signature of the active driving. For instance, a flux of photons has a different time-reversal signature than chemical driving [38]., and we focus on the case of chemical reactions for illustration purposes. The interactions between active particles are described by the free energy functional $F([\phi],a)$, which should not be confused with the Helmholtz free energy bounding $W_{\rm ext}$ in passive systems [1]. The system is maintained away from equilibrium by setting the difference of some chemical potentials $\Delta\mu$ constant; for instance, involving reactant and product molecules whose reaction fuels the particle activity [38].
Given that the total number of active particles is conserved, $\phi$ obeys a conservation law in terms of the diffusive current ${\boldsymbol{J}}$:
$\dot{\phi}+\boldsymbol{\nabla}\cdot\boldsymbol{J}=0\,,$ (3)
where $\dot{\phi}=\partial_{t}\phi$. We describe the stochastic coupling of $(\phi,n)$ through a linear relation between thermodynamic fluxes $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and forces $\mathbb{F}=(-\boldsymbol{\nabla}(\delta F/\delta\phi),\Delta\mu)$ [38]:
$-ν\mathbb{J}=\mathbb{L}\cdot\mathbb{F}+T^{1/2}\mathbb{\Lambda}+T\bbnu\,.$ (4)
where the noise term $\mathbb{\Lambda}$ is Gaussian with zero mean and correlations given by:
$\left\langle\mathbb{\Lambda}({\bf r},t)\mathbb{\Lambda}^{\intercal}({\bf r}^{\prime},t^{\prime})\right\rangle=2\mathbb{L}\delta({\bf r}-{\bf r}^{\prime})\delta(t-t^{\prime})\,.$ (5)
Here, refers to vector transpose, and we have set the Boltzmann constant to unity. Whenever $\mathbb{L}$ depends on $\phi$ and its gradients, the noise $\mathbb{\Lambda}$ is multiplicative. Then, interpreting Eq. (4) within the Stratonovitch convention [59, 60], the spurious drift $-ν\bbnu$ can be explicitly calculated for any given $\mathbb{L}$ (see Ref. [61], section VI).
For $d$ spatial dimensions, $\mathbb{J}$ and $\mathbb{F}$ are vectors in $d+1$ dimensions. The Onsager matrix $\mathbb{L}$ is symmetric and semi-positive definite:
$\mathbb{L}=\begin{pmatrix}\lambda_{\phi}\boldsymbol{1}_{d}&\boldsymbol{C}(\phi,\boldsymbol{\nabla}\phi)\\ \boldsymbol{C}^{\intercal}(\phi,\boldsymbol{\nabla}\phi)&\lambda_{n}\end{pmatrix}\,,$ (6)
where $(\lambda_{\phi},\lambda_{n})$ are mobility constants and $\boldsymbol{1}_{d}$ is the $d\times d$ identity matrix. Physically, $\delta F/\delta\phi$ represents the free-energy cost of adding a particle, while $\Delta\mu$ corresponds to the energy gain per fuel molecule during the chemical reaction. The diagonal terms in Eq. (6) indicate that the negative gradient of $\delta F/\delta\phi$ drives the diffusive current $\boldsymbol{J}$, while $\Delta\mu$ drives the chemical reaction rate $\dot{n}$. The off-diagonal terms in Eq. (6) couple the chemical reaction to the diffusive dynamics. In general, the coupling term $\boldsymbol{C}$ is a function of $\phi$ and its gradients. Through this coupling, $\Delta\mu$ indirectly drives $\phi$ to operate away from equilibrium.
Considering a given protocol $a(t)$ [Fig. 1], the expression of $Q_{\rm t}$ follows from the product between the thermodynamic fluxes and forces [51]:
$Q_{\rm t}=\left\langle\int_{0}^{\tau+\tau_{\rm r}}\int_{V}\mathbb{J}\cdot\mathbb{F}\,\mathrm{d}{\bf r}\mathrm{d}t\right\rangle\,,$ (7)
[…]
$W_{\text{ext}}=\int_{0}^{\tau}\dot{a}\left\langle\partial_{a}F\right\rangle\mathrm{d}t\,.$ (8)
The work produced by the chemical reaction over the entire process is given by
$W_{\text{act}}=\left\langle\int_{0}^{\tau+\tau_{\rm r}}\int_{V}\Delta\mu\dot{n}\,\mathrm{d}{\bf r}\mathrm{d}t\right\rangle\,,$ (9)
which can also be written as
$W_{\text{act}}=(\tau+\tau_{\rm r})P_{0}+\int_{0}^{\tau+\tau_{\rm r}}\langle P\rangle\mathrm{d}t\,,$ (10)
where the power $P_{0}=\lambda_{n}V\Delta\mu^{2}$ is a background contribution, independent of the coupling between the diffusive and chemical sectors of the dynamics, and
$P=\int_{V}p\,\mathrm{d}{\bf r}\,,\quad p=\frac{\Delta\mu}{\lambda_{\phi}}\boldsymbol{C}\cdot\left(\boldsymbol{J}-\Delta\mu\,\boldsymbol{C}\right)$ (11)
embodies how the $\phi$-dynamics affects the dissipation rate. It is worth noting that $W_{\text{act}}\geq 0$, as discussed in [38]. The expressions of heat and work [Eqs. (7-11)] straightforwardly extend to more complex field theories [38, 51], beyond the simple case of a single conserved scalar field [Eq. (3)].
The relation between the protocol heat $Q_{\rm p}$, the energy $E$, and the total work $W_{\rm t}$ is identical to that for $Q_{\rm t}$ in Eq. (1), with two key modifications: (i) $E(\tau+\tau_{\rm r})$ is replaced by $E(\tau)$, where $E(\tau)$ denotes a non-stationary average of $F$ at time $\tau$; and (ii) in the definition of $W_{\rm act}$ [Eq. (10)], the time $\tau+\tau_{\rm r}$ is replaced by $\tau$.
Importantly, the averages in Eqs. (8-11) are not stationary, and depend on the details of the protocol $a(t)$. Whenever $a$ is fixed ($\dot{a}=0$ and $W_{\rm ext}=0$), these averages reduce to their steady-state values, so that $W_{\rm act}$ scales like $\tau+\tau_{\rm r}$. In such a case, we have $E(0)=E(\tau)=E(\tau+\tau_{\rm r})$, and we deduce from Eq. (1) that both $(Q_{\rm p},Q_{\rm t})$ diverge with the protocol duration. This regime reproduces the scaling of slow protocols [Figs. 1(b-c)], and is a direct consequence of the permanent fuel consumption, which is present even when the parameter $a$ is kept constant.
Perturbative treatment for slow protocols
We offer a systematic approach to express the heats $(Q_{\rm p},Q_{\rm t})$ as functionals of the protocol $a(t)$ for a generic active field theory [Eqs. (3-6)]. Inspired by previous works [54, 4], we rely on a perturbative expansion around the quasistatic limit ($\Omega\tau\gg 1$, where $\Omega$ is the slowest relaxation frequency of the system (see Ref. [61], section IV.B), and assume that the protocol is smooth by neglecting any discontinuity in $a(t)$; see dynamical response theory (DRT) in Methods. Within DRT, we obtain an implicit expression of $(Q_{\rm p},Q_{\rm t})$ for an arbitrary protocol $a(t)$ as
$Q_{\rm t}$ $=(\tau+\tau_{\rm r})P_{0}+\tau_{\rm r}\langle P\rangle_{\text{s}}(a_{\tau})+B_{\rm t}+\int_{0}^{\tau}L(a,\dot{a})\mathrm{d}t\,,$ (12)
$Q_{\rm p}$ $=\tau P_{0}+B_{\rm t}+\dot{a}_{\tau}B_{\rm p}+\int_{0}^{\tau}L(a,\dot{a})\mathrm{d}t\,,$
where $\langle\cdot\rangle_{\text{s}}$ denotes a stationary average, and $(B_{\rm t},B_{\rm p})$ are boundary terms set by $(a_{0},a_{\tau})$ (see Ref. [61], section I). The Lagrangian $L$ is the only contribution depending on the full protocol:
$L(a,\dot{a})=m(a)\dot{a}^{2}+\langle P\rangle_{\text{s}}(a)\,.$ (13)
Within Lagrangian mechanics, $L$ maps into the Lagrangian of a particle with position $a$ and effective mass $m(a)$ subject to the potential $-\langle P\rangle_{\text{s}}(a)$. We relate $m(a)$ to some response functions (or, alternatively, some correlation functions [4]) measured at fixed $a$ (see Ref. [61], section I).
At large $\tau$, $L$ is dominated by the term $\langle P\rangle_{\text{s}}(a)$, so that $(Q_{\rm p},Q_{\rm t})$ scale like $\tau$. At short $\tau$, $L$ is dominated by the term $m(a)\dot{a}^{2}$, so that $(Q_{\rm p},Q_{\rm t})$ scale like $1/\tau$; although such a scaling deviates from the expectations for $Q_{\rm p}$ [Fig. 1], it can still help capture the location of any local minimum. In fact, the crossover between the DRT predictions at large and small $\tau$ always entails a minimum for $(Q_{\rm p},Q_{\rm t})$ at finite $\tau$. For a given $a(t)$, we deduce from Eqs. (12-13) the location of the optimal durations:
$\tau_{\rm p}^{2}=\frac{\dot{a}_{s=1}B_{\rm p}+\int_{0}^{1}\dot{a}^{2}m(a)\mathrm{d}s}{P_{0}+\int_{0}^{1}\langle P\rangle_{\text{s}}(a)\mathrm{d}s}\,,\quad\tau_{\rm t}^{2}=\frac{\int_{0}^{1}\dot{a}^{2}m(a)\mathrm{d}s}{P_{0}+\int_{0}^{1}\langle P\rangle_{\text{s}}(a)\mathrm{d}s}\,,$ (14)
where we have changed variable $s=t/\tau$, and $\dot{a}=da/ds$. Therefore, $Q_{\rm p}$ and $Q_{\rm t}$ are not minimum at the same duration $\tau$ in general. Again, let us emphasize that DRT predicts that $Q_{\rm p}$ always features a local minimum: we examine below in detail the conditions under which this prediction can actually fail. On the contrary, $Q_{\rm t}$ always features a minimum and DRT will give good predictions for it close to the quasistatic limit.
In short, the DRT predictions [Eqs. (12-14)] hold for a generic active field theory within our thermodynamically-consistent formulation [Eqs. (3-6)], yet they rely on assuming that the protocol is close to quasistatic ($\Omega\tau\gg 1$). To deploy DRT on a specific dynamics, the main challenge is to explicitly evaluate the Lagrangian in terms of $a(t)$. In what follows, we demonstrate how DRT can inform the optimization of protocols for a new model of active phase separation.
Field theory of active phase separation
We now introduce a case study, referred to as Chemical Model B (CMB), which corresponds to a specific active field theory for nonequilibrium phase separation. To this end, we consider the canonical $\phi^{4}$ free-energy functional [36]
$F([\phi],a)=\int_{V}\bigg{[}\frac{a(t)}{2}\phi^{2}+\frac{b}{4}\phi^{4}+\frac{\kappa}{2}(\boldsymbol{\nabla}\phi)^{2}\bigg{]}\mathrm{d}x\,,$ (15)
where $b,\kappa>0$. We focus on the case of one spatial dimension ($d=1$), for which the spurious drift $-ν\bbnu$ [Eq. (4)] vanishes [38]. Here, the parameter $a(t)$ is related to the second virial coefficient, as obtained from coarse-graining procedures [62, 63], and can be varied in various manners (e.g., controlling particle interactions by changing the pH of tunable colloidal systems [64]).
To leading order in powers of $\phi$ and its gradients, the simplest form of the coupling term $C$ [Eq. (6)] reads
$C=\gamma\partial_{x}\phi\,,$ (16)
where $\gamma$ is constant. The expression in Eq. (16) ensures that $C$ vanishes for a constant $\phi$, as expected for homogeneous isotropic systems. The corresponding dynamics of the active field [Eqs. (3-4)] follows as
$\dot{\phi}$ $+\partial_{x}J=0\,,$ (17)
$J$ $=-\lambda_{\phi}\partial_{x}\bigg{[}\bigg{(}a-\frac{\gamma\Delta\mu}{\lambda_{\phi}}+b\phi^{2}-\kappa\partial_{x}^{2}\bigg{)}\phi\bigg{]}+T^{1/2}\Lambda_{\phi}\,,$
where the noise term $\Lambda_{\phi}$ [Eqs. (5-6)] is Gaussian with zero mean and correlations given by
$\langle\Lambda_{\phi}(x,t)\Lambda_{\phi}(x^{\prime},t^{\prime})\rangle=2\lambda_{\phi}\delta(x-x^{\prime})\delta(t-t^{\prime})\,.$ (18)
Therefore, at the level of the $\phi$-dynamics, the coupling between chemical and active fields [Eq. (16)] simply amounts to taking $a\to\tilde{a}=a-\gamma\Delta\mu/\lambda_{\phi}$ in the free energy $F$ [Eq. (15)]. Then, the statistics of $\phi$ obeys equilibrium properties (Boltzmann steady-state, fluctuation-dissipation theorem, etc. [59]), although the coupled dynamics of $(\phi,n)$ operates away from equilibrium.
Figure 2: (a) Phase diagram of Chemical Model B (CMB) [Eqs. (15-18)] in terms of the free-energy parameter $\tilde{a}=a-\gamma\Delta\mu/\lambda_{\phi}$ and the average density $\phi_{0}=\frac{1}{V}\int_{V}\phi\mathrm{d}x$. The solid blue line ($\tilde{a}_{\text{1st}}=-b\phi_{0}^{2}$) separates homogeneous and phase-separated states with a critical point at the red dot. The color map indicates the value of $\langle P\rangle_{\text{s}}$ for $T>0$ [Eqs. (19-20)]. (b) The mean-field ($T=0$) dissipation rate $P$ changes across phase transitions when varying activity $\Delta\mu$, from $P=0$ (homogeneous) to $P<0$ (phase separation). Through the critical point ($\phi_{0}=0$), $P$ changes smoothly. For a discontinuous transition ($\phi_{0}=0.5$), $P$ changes more steeply as $V/\ell$ increases. Inset shows, for $\phi_{0}=0$ and in the phase-separated state, the spatial profiles of the density $\phi$ (red solid line) and local power $p$ (blue solid line). Parameters: $a=b=\kappa=\lambda_{\phi}=\gamma=1$, $V/\ell=64$, $\Delta\mu=3$, (a) $F(\phi_{0},a_{0})/T=50V/\ell$, (b) $T=0$.
The equilibrium mapping of the active field dynamics allows us to readily deduce the phase diagram [36]. For large enough $\Delta\mu$, the system undergoes a separation between high-density ($\phi>0$) and low-density ($\phi<0$) regions [Fig. 2(a)]. Unlike other models of active phase separation [39, 65, 51], the activity in CMB does not necessarily stem from microscopic self-propulsion. More generally, activity here results from a coupling of the system with a fuel reservoir; for instance, ATP molecules powering nonequilibrium phase separation in living cells [66, 67].
For a constant $a(t)$, the active work $W_{\rm act}$ quantifies the dissipation required to sustain steady-state activity, and follows from Eqs. (10-11) as
$W_{\rm act}=(\tau+\tau_{\rm r})(P_{0}+\langle P\rangle_{\text{s}})\,,$ (19)
where now
$P=\int_{V}p\,\mathrm{d}x\,,\quad p=\frac{\gamma\Delta\mu}{\lambda_{\phi}}(\partial_{x}\phi)\left(J-\gamma\Delta\mu\,\partial_{x}\phi\right)\,,$ (20)
and $J$ is given in Eq. (17). At the mean-field level ($T=0$), the phase-separated profile $\phi(x)$ is known analytically [36], from which we deduce $p(x)$ (see Ref. [61], section IV.A): it vanishes in the bulk phases, and is negative at the interface [Fig. 2(b) inset]. Therefore, the dissipation is maximal in the homogeneous state, where all the fuel energy is wasted in the thermostat, and is reduced for phase separation, where part of the fuel energy serves to shape the density profile [Fig. 2(b)]. In the homogeneous state, our analytical results (see Ref. [61], section IV.B) show that the finite-temperature value of $\langle P\rangle_{\text{s}}$ is negative, and its absolute value reduces closer to the phase boundary [Fig. 2(a)], while maintaining $W_{\text{act}}\geq 0$.
The thermodynamically consistent structure of our active field theory [Eqs. (3-6)] is built to properly account for all sources of dissipation coming from the diffusive and chemical sectors of the dynamics. For CMB [Eqs. (15-18)], the dissipation is locally reduced at interfaces, in line with similar results for another model of active phase separation [38]. As discussed below, the behavior of the steady-state dissipation across the phase diagram determines the shape of the optimal protocol.
Figure 3: (a) Protocol heat $Q_{\rm p}$ and (b) total heat $Q_{\rm t}$ as functions of the protocol duration $\Omega\tau$ for the linear protocol $a_{\rm lin}$ [Eq. (21)] at two levels of activity $\Delta\mu/T$. Markers refer to numerical simulations (code available in [68]), solid lines to analytical results, and dashed lines to predictions of DRT [Eqs. (12)]. Parameters: $a_{0}=1$, $a_{\tau}=10$, $b=\kappa=1$, $\phi_{0}=0.5$, $\lambda_{\phi}=\lambda_{n}=\gamma=1$, $V/\ell=64$, $F(\phi_{0},a_{0})/T=50V/\ell$.
Scalings of optimal protocol duration
For CMB dynamics [Eqs. (15-18)], we consider a protocol that evolves $a(t)$ within the homogeneous state. Without loss of generality, we take $a_{\tau}>a_{0}$ [Fig. 2(a)], so that the energy difference $\langle F(a_{\tau})-F(a_{0})\rangle$ and the heat $(Q_{\rm p},Q_{\rm t})$ are always positive. For simplicity, we first examine the case of the linear protocol:
$a_{\rm lin}(t)=a_{0}+(t/\tau)(a_{\tau}-a_{0})\,.$ (21)
Using numerical simulations (see Ref. [61], section IV.G), we evaluate the heat $(Q_{\rm p},Q_{\rm t})$ [Eqs. (1-2) and (8-11)] for two levels of activity $\Delta\mu/T=(0.1,0.5)$ [Fig. 3]. In both cases, $Q_{\rm t}$ features a global minimum at an intermediate value of the protocol duration $\tau$. In contrast, $Q_{\rm p}$ has a local minimum for $\Delta\mu/T=0.1$, whereas it increases monotonically for $\Delta\mu/T=0.5$. As previously discussed, these features support the fact that $Q_{\rm t}$ is generally the relevant heat to optimize, since the existence of a minimum is independent of the activity.
To obtain some analytical predictions, we focus on the regime of weak noise where $T$ is much smaller than the local energy scale $(\ell/V)F$, with $\ell$ being the grid spacing, which also corresponds to large system size. Expanding the active field $\phi$ around the mean-field solution $\phi_{0}=\frac{1}{V}\int_{V}\phi\mathrm{d}x$, the dynamics of each mode $\tilde{\phi}_{k}=\frac{1}{V}\int_{V}e^{ikx}(\phi-\phi_{0})\mathrm{d}x$ is decoupled and linear to leading order in $T$. Our corresponding analytical results for $(Q_{\rm p},Q_{\rm t})$ (see Ref. [61], section III.B) reproduce the curves sampled numerically [Fig. 3]. Alternatively, one can also evaluate $(Q_{\rm p},Q_{\rm t})$ using DRT in the weak-noise regime (see Ref. [61], section III): these predictions are in good agreement with the numerics at large $\Omega\tau$, yet show a discrepancy at small $\Omega\tau$ [Fig. 3], as expected. Moreover, DRT quantitatively captures the locations of $(\tau_{\rm t},\tau_{\rm p})$ only at a small activity ($\Delta\mu/T=0.1$).
We now analyze how the extremal durations $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$ [Fig. 1] scale with activity $\Delta\mu/T$ and system size $V/\ell$. In particular, our aim is to find the regimes where the DRT predictions are accurate. Based on leading-order expansions (see Ref. [61], section IV.D), we obtain the scalings $(B_{\rm p},m)={\cal O}(V^{2},\Delta\mu^{0})$ and $\left(P_{0},\langle P\rangle_{\text{s}}\right)={\cal O}(V,\Delta\mu^{2})$ at large $V$ and small $\Delta\mu$, from which we deduce the scalings of the minima $(\tau_{\rm p},\tau_{\rm t})={\cal O}(V^{1/2},\Delta\mu^{-1})$ using DRT predictions [Eq. (14)]. Our analytical results for $(\tau_{\rm p},\tau_{\rm t})$ are in good agreement with DRT for small $\Delta\mu/T$ and moderate $V/\ell$ [Figs. 4(a-b)], and the relative errors increase with $(\Delta\mu/T,V/\ell)$ [Figs. 4(c-d)]. Remarkably, the analytical results also suggest the scaling of the maximum $\tau_{\rm max}={\cal O}(V^{2},\Delta\mu^{0})$, from which it follows that the local minimum and maximum of $Q_{\rm p}$ [Fig. 1] always merge at large $(\Delta\mu/T,V/\ell)$ [Figs. 4(a-b)].
We reveal that DRT becomes less accurate as volume $V$ and activity $\Delta\mu$ increase: this result is consistent with the assumption that DRT breaks down when the relaxation time $\Omega^{-1}$, which decreases with $(V,\Delta\mu)$ (see Ref. [61], section IV.B), becomes comparable with $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$. In fact, our DRT scaling predictions are valid for an arbitrary protocol $a(t)$. The scaling $(\tau_{\rm p},\tau_{\rm t})\sim V^{1/2}$ is a consequence of the conservation of particle number: we find a different scaling for active theories of non-conserved fields (see Ref. [61], section V). In contrast, the scaling $(\tau_{\rm p},\tau_{\rm t})\sim\Delta\mu^{-1}$ is robust: it embodies the competition between internal activity and external perturbation, respectively regulating the small- and large-$\tau$ regimes, inherent to any active system [4].
Figure 4: Extremal protocol durations $(\tau_{\rm p},\tau_{\rm t},\tau_{\rm max})$ of the heat $(Q_{\rm p},Q_{\rm t})$ [Fig. 1] as functions of (a) system size $V/\ell$ and (b) activity $\Delta\mu/T$ for the linear protocol $a_{\rm lin}$ [Eq. (21)]. Solid and dashed lines respectively refer to analytical results and DRT predictions [Eq. (14)]. Dotted lines are guidelines for scaling behaviors. Beyond the shaded region, $\tau_{\rm p}$ and $\tau_{\rm max}$ coalesce such that $Q_{\rm p}$ becomes monotonic. (c-d) The relative errors $\varepsilon_{\rm x}=\tau_{\rm x}^{\rm DRT}/\tau_{\rm x}^{\rm anal}-1$, defined for ${\rm x}=({\rm p},{\rm t})$, compare analytical results with DRT predictions taken from (a-b), respectively. Parameters: $a_{0}=1$, $a_{\tau}=10$, $b=10$, $\kappa=1$, $\phi_{0}=0.5$, $\lambda_{\phi}=\lambda_{n}=\gamma=1$, $F(\phi_{0},a_{0})/T=50V/\ell$, (a,c) $\Delta\mu/T=0.02$, (b,d) $V/\ell=10$.
Crossover between optimal protocols
We continue by examining the DRT predictions for the optimal protocol $a_{\text{op}}(t)$ minimizing $(Q_{\rm p},Q_{\rm t})$ [Eq. (12)] at a given duration $\tau$. This amounts to optimizing the Lagrangian $L$ [Eq. (13)], since the rest of the terms in $Q_{\rm t}$ [Eq. (12)] are independent of the protocol shape $a(t)$. The optimization of $L$ is obtained explicitly by solving the corresponding Euler-Lagrange equation (see Ref. [61], section IV.E). Since $Q_{\rm p}$ and $Q_{\rm t}$ [Eq. (12)] have the same functional dependence on $a(t)$, we attribute them the same optimal protocol $a_{\rm op}(t)$. Again, we focus here on the CMB dynamics in the weak-noise regime [Eqs. (15-18)] with $a(t)$ evolving within the homogeneous state, see Fig. 5(a).
In general, the shape of the optimal protocol changes with $\tau$, that is $a_{\rm op}=a_{\rm op}(t/\tau,\tau)$. Interestingly, we observe a crossover between two types of optimal protocols: at small $\tau$, $a_{\text{op}}$ tends to the master curve $\alpha_{0}(t/\tau)$, which monotonically increases from $a_{0}$ to $a_{\tau}$; at large $\tau$, $a_{\text{op}}$ tends to the master curve $\alpha_{\infty}(t/\tau)$, which has a non-monotonic behavior [Fig. 5(a)]. We rationalize this crossover by analyzing how $L$ scales with $\tau$. Given that the optimal protocol takes the form $a_{\rm op}(t)=\alpha(t/\tau)$ at large or small $\tau$, the corresponding contribution to $(Q_{\rm p},Q_{\rm t})$ can be written as
$\int_{0}^{\tau}L(a_{\rm op},\dot{a}_{\rm op})\mathrm{d}t=\int_{0}^{1}\bigg{[}\frac{m(\alpha)\dot{\alpha}^{2}}{\tau}+\tau\langle P\rangle_{\text{s}}(\alpha)\bigg{]}\mathrm{d}s\,.$ (22)
The first term in Eq. (22) dominates at small $\tau$: it has the same functional form as for a passive system [54], and accordingly yields the monotonic protocol $\alpha_{0}(t/\tau)$. The second term in Eq. (22) dominates at large $\tau$: it is optimal when the protocol minimizes the steady-state power $\langle P\rangle_{\text{s}}$ [Eq. (20)]. For CMB, $\langle P\rangle_{\text{s}}$ is minimal at the phase boundary [Fig. 2(a)]. Therefore, the corresponding optimal protocol consists of setting $a(t)$ close to this boundary for as long as possible, leading to the master curve $\alpha_{\infty}(t/\tau)$ [Fig. 5(a)].
The predictions for the optimal heat $Q_{\rm t}$, which minimize heat for all protocol shapes and durations, follow by substituting $a_{\rm op}$ into Eq. (12). The agreement with numerical measurements breaks down at small $\Omega\tau$, as expected, yet the discrepancy remains negligible close to the (global) optimal duration $\tau_{\rm t}$ [Fig. 5(b)]. By matching the asymptotic behaviors of $Q_{\rm t}$ at large and small $\tau$, we approximate $\tau_{\rm t}$ [Eq. (14)] as
$\tau_{\rm t}^{2}\approx\frac{\int_{0}^{1}\dot{\alpha}_{0}^{2}m(\alpha_{0})\mathrm{d}s}{P_{0}+\int_{0}^{1}\langle P\rangle_{\text{s}}(\alpha_{\infty})\mathrm{d}s}\,.$ (23)
In practice, Eq. (23) reproduces the value estimated numerically (see Ref. [61], section IV.F for more details) for our specific parameters [Fig. 5(b)]: DRT predictions quantitatively capture the global minimum of $Q_{\rm t}$ whenever their regime of validity contains $\tau_{\rm t}$.
Therefore, we conclude that the crossover between optimal protocols is governed by the competition between the dissipation stemming from either external driving or internal activity, respectively dominant at small and large $\tau$. This crossover is expected to appear generically when controlling homogeneous states in a broad class of active systems, beyond the specific case of CMB (see Ref. [61], section IV.E).
Figure 5: (a) Optimal protocol $a_{\text{op}}$, minimizing $Q_{\rm t}$ for a given protocol duration $\tau$, as a function of time $t/\tau$. Colors refer to values of $\tau$, from $\Omega\tau\ll 1$ (dark blue, master curve $\alpha_{0}$) to $\Omega\tau\gg 1$ (dark red, master curve $\alpha_{\infty}$). Inset shows the protocols $(\alpha_{0},\alpha_{\infty})$ in the phase diagram [Fig. 2]. (b) Total heat $Q_{\rm t}$, optimized for a given $\tau$, as a function of protocol duration $\Omega\tau$. The solid line, with color code corresponding to protocols in (a), and markers respectively represent DRT predictions and numerical simulations. The dashed black lines are the asymptotic behaviors extracted from master curves in (a). The red horizontal dashed line highlights the (global) optimal protocol duration $\tau_{\rm t}$ [Eq. (23)]. Same parameters as in Fig. 3.
Discussion
In this work, we have examined how to quantify and optimize the heat dissipated by complex active systems under external perturbation. Specifically, for thermodynamically consistent field theories of active matter, we have distinguished two definitions of heat. The protocol heat $Q_{\rm p}$ is dissipated only during the protocol, and may feature a local minimum as a function of the protocol duration $\tau$ only under specific conditions. In contrast, the total heat $Q_{\rm t}$ is dissipated during both the protocol and its post-relaxation, and systematically features a global minimum. Therefore, our study goes beyond previous works which mostly focused on $Q_{\rm p}$ [15, 4]: we have here highlighted the essential role of post-relaxation in quantifying the energetics of active systems.
We have put forward a systematic framework, inspired by previous works [54, 4], to optimize the external perturbation using dynamical response theory. Other frameworks unrealistically assume that the operator has full control over the system (namely, by completely changing the rules of the dynamics) [69, 70]; this assumption is particularly useful to derive bounds on dissipation [71, 72, 73]. In contrast, our predictions apply here to the manipulation of some experimentally accessible parameters (namely, by changing the shape of a given free-energy). Remarkably, our framework allows one to evaluate: (i) the optimal duration of operation for a given protocol, (ii) the optimal protocol for a fixed duration, and (iii) the overall duration that gives minimal heat and the corresponding optimal protocol.
For the dynamics of a conserved scalar field that exhibits a nonequilibrium phase-separation, our scaling analysis predicts that $Q_{\rm p}$ does not feature any minimum at large activity $\Delta\mu$ and volume $V$. A similar analysis for non-conserved scalar fields yields a different scaling for $V$ (see Ref. [61], section V), so that $Q_{\rm p}$ may show a local minimum even at large $V$. These results pave the way towards a systematic scaling analysis to inform protocol optimization in a broader class of active field theories. For instance, our approach could be deployed in polar active theories of the flocking transition [40], for which a thermodynamically-consistent description has been recently proposed [49].
When driving homogeneous active systems, we have revealed a crossover between two families of optimal protocols. In the specific case of a conserved scalar field, we have shown that this crossover embodies the competition between the internal and external contributions to dissipation, stemming respectively from activity and perturbation. Remarkably, since this competition is inherent to a broad class of active systems [4], optimizing protocols between homogeneous states should generically yields a similar crossover. To go beyond our results, it remains to address protocols with multiple control parameters, for instance to build thermodynamic cycles [74, 75, 76]. Moreover, it would be interesting to examine how our framework can be extended to protocols crossing phase transitions.
Acknowledgements.
AS and TM acknowledge funding from the Israel Science Foundation (Grant No. 1356/22). ET acknowledges funding from EPSRC (Grant No. EP/W027194/1). EF acknowledges funding from the Luxembourg National Research Fund (FNR), grant reference 14389168. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Methods
Stochastic thermodynamics
Stochastic thermodynamics provides a way to calculate the average heat dissipated into the environment during the time $\tau+\tau_{\rm r}$ as [13, 60]:
$Q_{\rm t}=T\left\langle\ln\frac{\mathcal{P}_{+}(\tau+\tau_{\rm r})}{\mathcal{P}_{-}(\tau+\tau_{\rm r})}\right\rangle\,,$ (24)
where $\mathcal{P}_{+}$ and $\mathcal{P}_{-}$ are the probabilities of observing a given trajectory with the forward and the time-reversed dynamics, respectively. These probabilities take the form $\mathcal{P}_{\pm}(t)\sim e^{-A_{\pm}(t)}$ in terms of the dynamical action:
$A_{\pm}(t)=\frac{1}{4T}\int_{0}^{t}\int_{V}\left(\mathbb{L}\cdot\mathbb{F}\mp\mathbb{J}\right)\cdot\mathbb{L}^{-1}\cdot\left(\mathbb{L}\cdot\mathbb{F}\mp\mathbb{J}\right)\mathrm{d}{\bf r}\mathrm{d}t^{\prime}\,,$ (25)
where $\mathbb{L}^{-1}$ is the inverse of the Onsager matrix $\mathbb{L}$. The expression in Eq. (25) is given in the Stratonovitch convention [77, 78], where we have omitted some terms which are invariant under time-reversal. This connection between path probabilities and the average heat dissipation is quite general and extends even to systems in which the Einstein relation does not hold [50].
Substituting Eq. (25) into Eq. (24) yields the total heat $Q_{\rm t}$ in terms of the flux $\mathbb{J}$ and force $\mathbb{F}$ [Eq. (7)], which can also be written as
$Q_{\rm t}=\left\langle\int_{0}^{\tau+\tau_{\rm r}}\int_{V}\bigg{(}\Delta\mu\,\dot{n}-\frac{\delta F}{\delta\phi}\dot{\phi}\bigg{)}\mathrm{d}{\bf r}\mathrm{d}t\right\rangle\,,$ (26)
where we have used the definitions $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and $\mathbb{F}=(-\boldsymbol{\nabla}(\delta F/\delta\phi),\Delta\mu)$. A related expression has been derived for the heat in the absence of perturbation ($\dot{a}=0$) in terms of steady-state averages [38]. For a given protocol $a(t)$, the standard chain rule, valid within the Stratonovitch convention [59], yields
$\int_{V}\frac{\delta F}{\delta\phi}\dot{\phi}\,\mathrm{d}{\bf r}=\dot{F}-\dot{a}\partial_{a}F\,,$ (27)
from which we deduce
$Q_{\rm t}=W_{\text{ext}}+W_{\text{act}}+E(t=0)-E(t=\tau+\tau_{\rm r})\,,$ (28)
where we have used $E=\langle F\rangle$, along with the definitions of the external work $W_{\text{ext}}$ [Eq. (8)] and the active work $W_{\text{act}}$ [Eq. (9)]. Substituting the dynamics [Eqs. (4-6)] into the definition of $W_{\text{act}}$, we obtain
$W_{\rm act}=(\tau+\tau_{\rm r})P_{0}+\left\langle\int_{0}^{\tau+\tau_{\rm r}}(P+P_{\text{spu}})\mathrm{d}t\right\rangle\,,$ (29)
in terms of
$P_{\text{spu}}$ $=T^{1/2}\frac{\Delta\mu}{\lambda_{\phi}}\int_{V}(\lambda_{\phi}\Lambda_{n}-\boldsymbol{C}\cdot\boldsymbol{\Lambda}_{\phi})\mathrm{d}{\bf r}$ (30)
$\quad+T\frac{\Delta\mu}{\lambda_{\phi}}\int_{V}(\lambda_{\phi}\nu_{n}-\boldsymbol{C}\cdot\boldsymbol{\nu}_{\phi})\mathrm{d}{\bf r}\,.$
The stochastic integrals in Eq. (29) are interpreted within the Stratonovitch convention. Consequently, we demonstrate in Ref. [61], section VI that $\langle\int P_{\text{spu}}\mathrm{d}t\rangle=0$, so Eq. (10) indeed corresponds to $W_{\rm act}$. In fact, the spurious drift $-ν\bbnu$ ensures thermodynamic consistency of the dynamics [Eq. (4)], but the active work [Eq. (10)] does not explicitly depend on $-ν\bbnu$, as expected for noise-averaged quantities.
Dynamical response theory
Evaluating the expressions for heat [Eq. (7)] and work [Eqs. (8-10)] requires computing the non-steady-state average $\left\langle O\right\rangle(t)$ of some observables $O$ at an arbitrary time $t$. In general, $\left\langle O\right\rangle(t)$ depends on the entire history of the protocol $a(t^{\prime})$ for $t^{\prime}\in[0,t]$. Inspired by previous studies [54, 4], in the regime of slow protocols ($\Omega\tau\ll 1$), we demonstrate how to express $\left\langle O\right\rangle(t)$ in terms of some specific response functions, or equivalently in terms of unperturbed correlation functions. Importantly, our approach assumes that the change in $a$ is slow compared with the system relaxation time $\Omega^{-1}$, yet the overall change $a(\tau)-a(0)$ can be large.
Let us split the protocol $a(t)=\sum_{n}\Delta a(t_{n},t_{n-1})$ as a series of discrete jumps $\Delta a(t_{n},t_{n-1})=a(t_{n})-a(t_{n-1})$ with durations $\delta t\ll\tau$, so that $t_{n}=n\delta t$ [79]. We approximate $\left\langle O\right\rangle(t_{n})$ close to $\left\langle O\right\rangle(t_{n-1})$ as [80]
$\left\langle O\right\rangle(t_{n})$ $\approx\left\langle O\right\rangle(t_{n-1})$ (31)
$\quad+\int^{t_{n}}_{t_{n-1}}\mathrm{d}t^{\prime}\,\Delta a(t^{\prime},t_{n})\bigg{[}R_{O}^{(1)}(t_{n}-t^{\prime})$
$\quad+\int^{t_{n}}_{t_{n-1}}\mathrm{d}t^{\prime\prime}\,\Delta a(t^{\prime},t_{n})R_{O}^{(2)}(t_{n}-t^{\prime},t_{n}-t^{\prime\prime})\bigg{]}\,,$
where $R_{O}^{(1)}$ and $R_{O}^{(2)}$ are the first- and second-order response functions, given respectively by
$R^{(1)}_{O}(a(t);t-t_{1})$ $=\left.\frac{\delta\left\langle O(t)\right\rangle}{\delta a(t_{1})}\right|_{a(t_{1})\to a(t)},$ (32a)
$R^{(2)}_{O}(a(t);t-t_{1},t-t_{2})$ $=\left.\frac{\delta^{2}\left\langle O(t)\right\rangle}{\delta a(t_{2})\delta a(t_{1})}\right|_{\begin{subarray}{c}a(t_{1})\to a(t)\\ a(t_{2})\to a(t)\end{subarray}}.$ (32b)
Provided that the system relaxes faster than the rate of change of $a$ within each timestep ($\Omega\delta t\gg 1$), we approximate $\langle O\rangle(t_{n-1})\approx\langle O\rangle_{\rm ss}(a(t_{n-1}))$ as a steady-state average, and we extend the lower limits of integration in Eq. (31) as $t_{n-1}\to-\infty$. Additionally, assuming that the protocol $a(t)$ is smooth, we expand
$\Delta a(t^{\prime},t_{n})\approx(t^{\prime}-t_{n})\dot{a}(t_{n})+(t^{\prime}-t_{n})^{2}\ddot{a}(t_{n})/2\,.$ (33)
Substituting this expansion in the response framework [Eq. (31)], we deduce
$\left\langle O\right\rangle(t)$ $\approx\left\langle O\right\rangle_{\text{s}}\bigl{(}a(t)\bigr{)}+\dot{a}(t)\zeta_{O}^{(1,1)}\bigl{(}a(t)\bigr{)}$ (34)
$\quad+\ddot{a}(t)\zeta_{O}^{(2,1)}\bigl{(}a(t)\bigr{)}+\dot{a}^{2}(t)\zeta_{O}^{(2,2)}\bigl{(}a(t)\bigr{)}\,,$
where we have taken $t\equiv t_{n}\approx t_{n-1}$, and introduced
$\zeta_{O}^{(1,1)}$ $=-\int_{0}^{\infty}t^{\prime}R_{O}^{(1)}(t^{\prime})\mathrm{d}t^{\prime}\,,\quad\zeta_{O}^{(2,1)}=\int_{0}^{\infty}\frac{t^{\prime 2}}{2}R_{O}^{(1)}(t^{\prime})\mathrm{d}t^{\prime}\,,$ (35)
$\zeta_{O}^{(2,2)}$ $=\iint_{0}^{\infty}\frac{t^{\prime}t^{\prime\prime}}{2}R_{O}^{(2)}(t^{\prime},t^{\prime\prime})\mathrm{d}t^{\prime}\mathrm{d}t^{\prime\prime}\,.$
Note that the observable-specific quantities in the rhs of Eq. (34), namely $\left\langle O\right\rangle_{\text{s}}$ and $\zeta_{O}^{(m,n)}$, depend only on $a(t)$ and not on its derivatives $(\dot{a},\ddot{a})$. Finally, substituting Eq. (34) into the expression of heat [Eq. (7)] yields the decomposition given in Eq. (12); see Ref. [61], section I for details.